OpenStudy (anonymous):

Multiply. (5√2+3√3)^2

OpenStudy (anonymous):

50+27+2*5sqrt(2) * 3sqrt(3) = 77 + 30sqrt(6). This is the answer.

OpenStudy (theviper):

\[=(5\sqrt{2})^2+(3\sqrt{3})^2+2(5\sqrt{2})(3\sqrt{3})\]

OpenStudy (theviper):

Can u do it now?

OpenStudy (theviper):

\[=50+27+30\sqrt{6}\]

OpenStudy (theviper):

i.e,\[77+30\sqrt{6}\]